OpenStudy (amber15):

A wheel made 352 revolutions at 4 3/5 revolution per minute how many minutes did it rotate

OpenStudy (anonymous):

the rpm is 23/5 now (rpm)(total minutes)=352 rev therefore total minutes = 352*5/23 total minutes = 76.5 or full revolutions 76

OpenStudy (amber15):

thank you

OpenStudy (anonymous):

atleast give me a medal

OpenStudy (anonymous):

let the time be t so \[\frac{ 352 }{ t } = 4\frac{ 3 }{ 5 } = \frac{ 23 }{ 5 }\] so on simplification t= 76.52